When the cold junction of a thermocouple thermometer is at 5 C and the hot junction at 110 C, the millivoltmeter shows a deflection of 3.5 mV towards the left. The hot junction is now put in an unknown temperature while maintaining the cold junction at 5 C. The millivoltmeter now shows a deflection of 4.7 mV towards the left. Calculate the unknown temperature.


1
Expert's answer
2021-09-07T23:53:38-0400

"E = A * \Delta T" Eq. 1

Where A is a constant of the thermocouple material, E is the emf produced due to temp. difference between hot and cold junction.


Given T0 = 5

T1 = 110 and EMF = 3.5 mV

Therefore, 3.5 = A * (110-5)

Or "A = \frac{3.5}{105} = 0.033"


Now Given, E = 4.7 mV

T0 = 5

T1=?

Using eq. 1, we get,

4.7 = A * (T1 - 5)

4.7 = 0.033 T1 - 0.167

Or T1 =

"T1 = \frac{4.7 + 0.167}{0.033} = 146.15 deg . C"


Therefore, the hot junction temp. = 146 deg C